OpenStudy (anonymous):

find the variation constant and an equation of variation where y varies directly x and y = 32 when x =4

OpenStudy (anonymous):

xy = 128

OpenStudy (anonymous):

constant is 128 equation is x = 128/y

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

the variation constant is k = and the equation ovariayion is y =

OpenStudy (anonymous):

is x^14 *X^0=x^14